OpenStudy (anonymous):

Simplify. 2a + 2 / a^2 - 1

OpenStudy (anonymous):

the numerator 2a+2 simplifies to 2(a+1) the denominator simplifies to (a+1)(a-1) and now we have 2(a+1)/(a+1)(a-1) = 2/(a-1) :)

OpenStudy (anonymous):

what about this one ? What is the undefined value of 3b^2 + 13b + 4 / b + 4

OpenStudy (anonymous):

when b=-4

OpenStudy (anonymous):

because the the undefined values or discontinuity points are when the denominator is equal to zero so b+4=0 and that is b = -4